OpenStudy (anonymous):

How do you solve sin^2x+cosx=2

OpenStudy (anonymous):

So far I have replaced sin^2x with 1-cos^2x

OpenStudy (anonymous):

rewrite \(\sin^2(x)\) as \(1-\cos^2(x)\) so you will have nothing but cosines then solve the quadratic equation in cosine then solve for \(x\)

OpenStudy (anonymous):

Right now I have cos^2x+cosx+1=0 and I tried to complete the square but it didn't work out

OpenStudy (anonymous):

-cosx actually

OpenStudy (anonymous):

\[1-\cos^2(x)+\cos(x)=2\] \[\cos^2(x)-\cos(x)+1=0\]

OpenStudy (anonymous):

Yeah that's what I have

OpenStudy (anonymous):

no real solutions for sure

OpenStudy (anonymous):

i.e. \(x^2-x+1=0\) has no real solutions, so either there is a typo in the problem or they weren't paying attention when they wrote it

OpenStudy (anonymous):

Ok thanks because I got a negative under a radical :)